OpenStudy (anonymous):

find the unknown coordinate so the line through the points has the given slope. (-7,-4), (x,0) ; slope is 1/3

OpenStudy (anonymous):

y-(-4)=1/3(x-(-7)) y+4=1/3x+7/3 y=1/3x-5/3 so at x=0 y=-5/3

OpenStudy (anonymous):

\[y1-y2 \div x1-x2=slope\] \[-4-0/-7-x = 1/3\] x=5

OpenStudy (anonymous):

\[m=\frac{y2-y1}{x2-x1}\]Plug in what you know. Solve for x.

OpenStudy (anonymous):

5